OpenStudy (anonymous):

use the method of substitution to solve linear equations 7a=3-b and 7a=7-b

OpenStudy (shane_b):

Are you sure those equations are typed in correctly?

OpenStudy (anonymous):

Yes it Says 7a=7-b And 7a=3-b

OpenStudy (anonymous):

Use The Method Substitution to solve the system of linear equations

OpenStudy (shane_b):

Does it make sense that 7a can equal 7-b and 3-b?

OpenStudy (anonymous):

And How Did You Get That ?

OpenStudy (shane_b):

From your question, you're stating that:\[\large 7a=7-b=3-b\] which is impossible

OpenStudy (anonymous):

No Its Two Of Them Which Is 7a=3-b And The Other One Is 7a=7-b

OpenStudy (shane_b):

I understand there are two equations. However, they are both stating that "7a="...but the right sides could never be equal. \[3-b \neq 7-b\]

OpenStudy (shane_b):

Maybe I'm not explaining it well enough :) If 7a=7-b (first equation) then 7a cannot also be equal to 3-b (second equation).

OpenStudy (anonymous):

So The Answer I Will Pick Will Be No Solution

OpenStudy (shane_b):

There are no values of a or b that would make both equations be valid. YES!!

OpenStudy (shane_b):

Do you at least understand why there can be no solution?

OpenStudy (anonymous):

Yes I Understand ,Because If 7a=7-b (first equation) then 7a cannot also be equal to 3-b (second equation).

OpenStudy (anonymous):

use the eliminatiuon method to solve the system of equations a+8b=1 a+4b=-3

OpenStudy (anonymous):

No I ODnt HAve A Clue With This

OpenStudy (shane_b):

No problem...I'll walk you through this one. Using the elimination method you're looking to cancel things out when you add the equations to each other. Looking at both equations I can see that by multiplying the entire second equation by -1 (which doesn't change the equation) will allow me to add both equations together and eliminate the "a" variable. It would look like this: a+8b=1 -a-4b=3 Add those two equations to get: 4b=4 Therefore, b=1. Now that you know b=1, you can substitute that back into either of the original equations and solve for a.

OpenStudy (anonymous):

Okai and how will i do that i understand this part

OpenStudy (shane_b):

So if you understand that b=1, substitute it back into one of the other equations and solve for a as follows: Original equation: a+8b=1 After substituting b=1, that becomes: a+8(1)=1 Simplifying: a+8=1 Subtract 8 from both sides to get: a=-7 Now you know the solution: a=-7 and b=1

OpenStudy (anonymous):

Thank You So Much Now I totally Understand !

OpenStudy (anonymous):

Do You KNoww How To Solve Linear Equations By Graphing ?

OpenStudy (anonymous):

2x-y=0 5x+y=7

OpenStudy (shane_b):

Yes, put each equation in terms of y and then graph them. Are you allowed to use a graphing calculator?

OpenStudy (anonymous):

so what wiull i be graphing ?

OpenStudy (shane_b):

Ok, rearranging each equation in term of y you get:\[y=2x\]\[y=5x-7\] Graphing Calculator Method (TI-83/99): Click on the "Y=" button and insert both equations as Y1 and Y2. Then just hit the Graph button. By hand: Doing it by hand requires to you create a table of values. Pick a range for x, say -5 to 5, and then calculate the y value using each equation above. From there you'll have x,y values to plot. Since both are linear equations, you can simply connect the dots for each equations plotted points to get a full graph. Note that you can compare your graphs with this one to see if they look correct: http://www.wolframalpha.com/input/?i=graph%3A+y%3D2x%2C+y%3D5x-7